OpenStudy (anonymous):

please help me, i got final exam tomorrow integrate : [(2+arctan 3x)^(1/2) / 1+9x^2

OpenStudy (anonymous):

\[\int\limits_{}^{}(2+\tan ^{-1}3x)^{1/2}\div(1+9x ^{2})\]

OpenStudy (amistre64):

if u=2+tan^-1(3x) du = 3x/(1+9x^2) dx

OpenStudy (amistre64):

make that 3/(1+9x^2)

OpenStudy (amistre64):

[S] u^(1/2)/3 du

OpenStudy (amistre64):

2(2+tan^-1(3x)) --------------- +C ?? 9

OpenStudy (anonymous):

tq very much...^o^...are u a lecturer or something??

OpenStudy (amistre64):

no, just waiting for the sumer semester to start up :)

OpenStudy (amistre64):

i left out a bit there.... the ^(3/2)

OpenStudy (amistre64):

2(2+tan^-1(3x))^(3/2) ------------------- +C ?? 9 \[\frac{2 (2+\tan^{-1}(3x))^{3/2}}{9} +C\]

OpenStudy (anonymous):

sorry that was a wrong post for a different question